OpenStudy (dumbsearch2):

A local market wants to mix some $7 apples with some $8 oranges. (These cost are per pound.) How many pounds of each item should they mix to obtain a 25 pound mixture that sells for $7.52 per pound?

OpenStudy (dumbsearch2):

Thanks, guys in advance! :)

OpenStudy (anonymous):

\[\frac{7x+8y}{25}=7.52\] \[x+y=25\]

OpenStudy (anonymous):

Use the substitution method to solve this system/simultaneous equation.

OpenStudy (dumbsearch2):

So the answer is 7.62 and 25? @Azteck

OpenStudy (anonymous):

No... what are you talking about?

OpenStudy (anonymous):

do they add up to 25? you're meant to find x and y using the two equations above.

OpenStudy (anonymous):

And I told you to use the substitution method. Have you even done simultaneous equations or system of equations?

OpenStudy (dumbsearch2):

Would the elimination method work in this case?? @Azteck

OpenStudy (anonymous):

Yes, but would you like to waste time instead of finding the most efficient way in solving this question?

OpenStudy (dumbsearch2):

Ah no probably not. But @Azteck, I'm better at elimination. :)

OpenStudy (anonymous):

Being better at something doesn't mean you will solve it faster. Looking at that, you're meant to make a logical decision as to what way can get you the answer the fastest. In the second equation, all you have to do is move y or x to the RHS and you will have x or y as the subject.

OpenStudy (dumbsearch2):

So the answer would 12 apples and 13 oranges? Not quite sure about this, and thanks a lot for your help! :) @Azteck

OpenStudy (anonymous):

Correct.